OpenStudy (anonymous):

Prove the identity:

OpenStudy (anonymous):

\[\tan^2\theta-\sin^2\theta=\tan^2 \sin^2 \theta \]

OpenStudy (abmon98):

\[Tan^2\theta=Sin^2\theta/Cos^2\] \[Sin^2\theta/Cos^2-Sin^2\theta\] \[\sin ^2\theta-Cos^2\theta*Sin^2\theta/Cos^2\] \[Take the Sin^2\theta as Common factor\] \[Sin^2\theta(1-Cos^2\theta/Cos^2)\] \[ Sin^2 \theta*Sin^2\theta/Cos^2\] \[Tan^2\theta*Sin^2\theta \]

OpenStudy (anonymous):

Thanks!

OpenStudy (abmon98):

Your welcome :)

OpenStudy (anonymous):

I see that it's just mostly algebraic work, besides memorizing and applying the right identities. Hopefully I can do the rest of my problem sheet :)

OpenStudy (abmon98):

Good to hear that @casualstudent